OpenStudy (anonymous):

A pair of dice is tossed: what is the probability that the sum is 5? What is the probability that at least one dice shows a 2?

OpenStudy (anonymous):

for the first one i got 1/9

OpenStudy (anonymous):

\[\Omega=6^2=36\]

OpenStudy (anonymous):

there are 4 ways that you can get 5: 4+1 1+4 3+2 2+3

OpenStudy (anonymous):

so 4/36=1/9

OpenStudy (experimentx):

is answer 9/36 ??

OpenStudy (anonymous):

for part 1 or 2?

OpenStudy (anonymous):

i'm not sure what the answer is

OpenStudy (anonymous):

part 1 is 4/36

OpenStudy (experimentx):

i must be out of my mind ... i was mixing 1 and 2.

OpenStudy (experimentx):

II) 1/6*5/6 + 5/6*1/6 + 1/6*1/6

OpenStudy (anonymous):

2) should be 11/36

OpenStudy (anonymous):

ok i got number 1 right. how did u get 11/36?

OpenStudy (anonymous):